A 10 kg object is suspended from a spring causes the spring to lengthen by 5 cm down. The object is subjected to a force of 2cos (t / 4) N (newtons), and it moves in an environment where the viscosity force is 1N when the speed is 6 cm / s. If the object is set in motion from its position of equilibrium with an initial speed of 2 cm / s, formulate the initial value problem that describes the movement of the object.

1. Find the solution to this problem.

2. Determine the transient and permanent parts of the solution.

3. If the external force is replaced by a force of 4cosωt of frequency ω, find the value of ω for which the amplitude of the forced response is maximum.
Answer #1

forced damped vibrations mu + yutku = f(+) General Case m= 10 kg k=? f (t) = 2 cos (4A) lets find spring constant k ! 10 kPage : so Up (4) 1.02 cos(lt) 4 10-3 UdGADE for ų (A) we need to no External force is there solve the equation as if lou ofCIASSMAte Dato: Page : = 2 cm/s or 0.02 d ucts or ult) for to at 0.02 = 0.835G - 1462 G= - 1.02 X 10-3 C2 = 1.49810-3 Ult) So